OpenStudy (erinweeks):

Determine whether the relation is a function. {(-7, -7), (-7, -8), (-1, 4), (6, 5), (10, -1)}

OpenStudy (erinweeks):

@jim_thompson5910 can you help?

OpenStudy (erinweeks):

@timo86m can you help?

OpenStudy (erinweeks):

I have about 3 questions i need help with

OpenStudy (erinweeks):

am i supposed to graph it myself?

OpenStudy (anonymous):

dont plot

OpenStudy (anonymous):

well if you a visual learner plot.

OpenStudy (erinweeks):

Its not a function i plotted some of them but to me i kind of determined it wasnt

OpenStudy (anonymous):



OpenStudy (anonymous):

if you get something like that then it is not a function. what it means if you get to ys for one x

OpenStudy (anonymous):

like say that x is one. so if you get like 1,2 and 1,3 or really 1,a where a is anynumber.

OpenStudy (zzr0ck3r):

if at any time you have a x value that belongs to different y values it is not a function you have (7,-7) and (-7,-8) thus not a function

OpenStudy (erinweeks):

so i was correct its not.

OpenStudy (zzr0ck3r):

(-7,-7) and (-7,-8) thus not a function*

OpenStudy (anonymous):

right away you can tell that happens cuz x at -7 gets 2 y values -7 and - 8 (-7, -7), (-7, -8)

OpenStudy (erinweeks):

thats what i thought because -7 & -7 doesnt work thats how i determined it..

OpenStudy (erinweeks):

Use the graph of y = f(x) to graph the given function g. g(x) = -2f(x)

OpenStudy (erinweeks):

A.

OpenStudy (erinweeks):

B.

OpenStudy (erinweeks):

C.

OpenStudy (erinweeks):

D.

OpenStudy (anonymous):

another way to think of it is if you grap a pencil and hold it vertically to the graph and move it left to right and your pencil doesn't touh twice 2 points then it is a function.

OpenStudy (anonymous):

the rule -2f(x) means flip it vertically and i think stretch vertically by factor of 2.

OpenStudy (erinweeks):

im confused..

OpenStudy (zzr0ck3r):

I hate to get you back to the last problem but I don't think you understood why "thats what i thought because -7 & -7 doesnt work thats how i determined it.." what does not work about (-7,-7)?

OpenStudy (erinweeks):

because they are both negatives?

OpenStudy (zzr0ck3r):

its not the point(-7,-7) that makes this thing not a function its the two points (-7,-7) and (-7,-8)

OpenStudy (zzr0ck3r):

no that has nothing to do with it

OpenStudy (zzr0ck3r):

do you see how on those two points(-7,-7) and (-7,-8) they share a x values but the y values are different?

OpenStudy (anonymous):

noo erin Every x must have only one ploted point.

OpenStudy (erinweeks):

oh okayy i see what your saying now. I got it thanks !

OpenStudy (anonymous):

as for secont q the rule is the original graph is flipped up and down. ANd shrinked by a factor of 2 i think.

OpenStudy (erinweeks):

@timo86m can you help me finish my second problem and i need help with one more thing can you go to wolfram and graph \[y = -\frac{ 1 }{ 2 }x-6\] and copy the url and post it on here because im not grasping the concept of how to do it

OpenStudy (erinweeks):

So for the second question I think I can elimante C and D.

OpenStudy (anonymous):

 NOtice how it intercepts the y axis at -6 y=mx+b = -1/2 x -6 m b b is where you start at the y axis. -1/2 means you go down 1 and over 2.. then draw a line from the -6 to that point the down one over 2 and connect with a line.

OpenStudy (anonymous):

m=rise / run rise over run. Or if it is negative downward/ run.

OpenStudy (erinweeks):

the graph you just drew is that for the second question or my 3rd one?

OpenStudy (anonymous):

for the y= = -1/2 x -6

OpenStudy (erinweeks):

So it would look like this

OpenStudy (erinweeks):

& can you still help me with my second question??

OpenStudy (anonymous):

Yes that is how it looks like. You begin by putting a point at y axis at -6 then go down 1 over 2 cuzz your m is -1/2

OpenStudy (erinweeks):

okay so i got the right graph? & @timo86m can you help me with the second question?

OpenStudy (anonymous):

yes i can :)

OpenStudy (anonymous):

btw which is the f(x) graph the dotted line or solid line? does it say on your paper or online class or something?

OpenStudy (erinweeks):

Hey i already figured that one out. can you help me with one more. Find the inverse of the one-to-one function.

OpenStudy (erinweeks):

\[f(x)=\frac{ 2x+5 }{ 7 }\]

OpenStudy (anonymous):

inverse just means solve for x :)

OpenStudy (erinweeks):

A. \[f ^{-1}(x) = \frac{ 7 }{ 2x-5 }\] B. \[f ^{-1}(x)=\frac{ 7x+5 }{ 2 }\] C. \[f ^{-1}(x)=\frac{ 7x-5 }{ 2 }\]

OpenStudy (erinweeks):

im not understanding how to solve it

OpenStudy (anonymous):

y=f(x)=(2x+5)/7 y=(2x+5)/7 solve for x 7y=2x+5 multiply both side by 7 7y-5=2x subtrac 5 (7y-5)/2=x sway variables y=(7x-5)/2

OpenStudy (anonymous):

treat y as f^-1(x)